# Automated discovery and optimization of autonomous quantum error correction codes for a general open quantum system

### v1

@@ II Formulation of the autonomous QEC code design problem as an optimization problem
We start the search by making an initial guess for the AQEC code parameters. In most calculations, we make a random guess. The code space is chosen by taking two orthogonal but otherwise random vectors in the full Hilbert space. The matrices $\hat{b}$ and $\hat{O}$ are filled with complex numbers, such that each real or imaginary part of each matrix element is chosen from a uniform distribution in the range $[-0.5,0.5]$. One constraint is that $\hat{O}$ must be Hermitian. We therefore randomly generate the matrix elements above the diagonal and take their complex conjugates to complete the matrix. The diagonal matrix elements are kept at zero in our calculations.
@@
Taking the best available code space, we consider that any neighboring basis can be obtained from the current basis using the update unitary operator
@@
where $\lambda_{m}$ are (small) coefficients and $G_{m}$ are the generators of unitary operators in the full Hilbert space. We generate a possible basis for every generator, using the small infinitesimal coefficient $10^{-8}$ and construct an array for the fidelity for the different possible updates. This array allows us to calculate the gradient to maximize the fidelity improvement. More specifically, we obtain a single generator that contains the optimal amount of the different $G_{m}$ matrices. We can update the code words by applying to them the matrix $U_{\rm update}$ with the optimal generator multiplied by an update step size, which we set initially at $10^{-2}$. We keep moving along the direction of the gradient until the fidelity reaches a maximum and starts decreasing. When we identify that we have overshot the fidelity peak, we start zooming in on the peak region until we can identify the peak location to a precision of $10^{-8}$ in the update step size, i.e. until an update step of $10^{-8}$ in either direction decreases the fidelity rather than increase it. We update the code words to those that correspond to the point of maximum fidelity.
@@
For the controlled decay matrix $\hat{b}$, since it can be any $n\times n$ matrix, we take $2n^{2}$ possible updates for all the real and imaginary parts of all matrix elements, each time adding a small increment to one of the matrix elements and calculating the fidelity with the small update. We use the increment size $10^{-6}$. Once we have the $2n^{2}$ values of fidelity for all the possible updates, we calculate the gradient and determine the optimal direction for updating $\hat{b}$. We then move in the optimal update direction until we reach a maximum in the fidelity. We note here that although we set a maximum for the allowed update in a single iteration ($10^{4}$ to any real or imaginary part of any matrix element), this condition was never needed by our automated search algorithm. Including the condition was motivated by the fact that the optimal matrix $\hat{b}$ will have some infinitely large matrix elements, as we will discuss in Sec. III. Whenever a peak in the fidelity was encountered during the update process, which is what happened in all iterations of all calculations, we start zooming in on the peak region until we can identify the maximum point location to within $10^{-8}$. A similar procedure is used to update the matrix $\hat{O}$, incorporating the constraint that the operator must be Hermitian.
@@
A natural question to consider here is whether the gradient-based approach is suitable for the problem of AQEC search, or whether this problem is prone to suffering from the presence of local optima that hinder the search for the optimal code. Our use of this approach was inspired by the results of Refs. Yamamoto2005 ; Yamamoto2007 ; Kosut , which alternated between optimizing the code space and the recovery operators in the conventional, measure-then-correct QEC paradigm. Those studies demonstrated that the search does not contain local minima, such that the gradient-based search and the alternation between updating different components of the code are good elements in a QEC code discovery algorithm. Our results in Sec. III below show some indications that our search sometimes gets trapped in local optima and therefore fails to find optimal solutions. The authors of Ref. Wang made a similar observation and argued that the use of the average fidelity as an optimization cost function results in the appearance of local optima.
@@ III Results
III.1 Reference point: physical qubit with relaxation
To assess the AQEC protocols below, we start by establishing the reference point against which most of the protocols analyzed below should be evaluated.
@@ III.1 Reference point: physical qubit with relaxation
The fidelity formula in Eq. (6) can then be obtained by examining how much of each matrix element survives the relaxation dynamics. Concretely, we ignore the term $(1-e^{-\gamma t})\rho_{11}$ in the 00 matrix element, replace every $\rho_{ij}$ by 1, take the sum of all four matrix elements and divide by 4:
@@
One fidelity value that will serve as an important reference point below is the one obtained by setting $\gamma t=1$, because we use this setting in our code optimization calculations. With this setting, the fidelity under free evolution, i.e. when no error-correction action is taken, is $F_{0}=0.64523519$.
@@
III.2 Qudit with uniform decay rates
Figure 1: Fidelity and code space dependence on optimization iteration number $k$ for 4-level system with uniform decay rates (Sec. III.B) when we fix the matrices $\hat{b}$ and $\hat{O}$ at their optimal values for the 13 AQEC code described in the main text. The algorithm reaches an extremely good code basis within the first 10 iterations. The dotted line in Panel (a) represents the fidelity value $1-F=1.5\times 10^{-6}$, which is the expected value for the optimal basis.
Figure 2: Infidelity ($1-F$) and absolute values of the real and imaginary parts of the induced decay matrix elements $b_{i,j}$ as functions of optimization iteration number $k$ for 4-level system with uniform decay rates when we fix the code space and matrix $\hat{O}$ at their optimal values in the 13 code. The fidelity reaches $F\approx 0.9997$ after $10^{5}$ iterations. The finite slope at $k=10^{5}$ in Panel (a) suggests that further improvement in the fidelity can be achieved with further optimization. In Panel (b), the two lines that keep increasing up to $k=10^{5}$ correspond to the real and imaginary parts of the same matrix elements. Depending on the random initial guess used in the search, in some computational runs the real part is higher, while in other runs the imaginary part is higher. Each one of these two lines is itself two lines that are extremely close to each other and are indistinguishable at the scale of this plot. Each such indistinguishable pair correspond to the two matrix elements $b_{1,0}$ and $b_{3,2}$, which should indeed be equal in the optimal code.
Figure 3: Infidelity ($1-F$) and absolute values of the real and imaginary parts of the coherent coupling matrix elements $O_{i,j}$ as functions of optimization iteration number $k$ for 4-level system with uniform decay rates when we fix the code space at the optimal choice $\left\{\left|1\right\rangle,\left|3\right\rangle\right\}$ and set $\hat{b}$ as in Eq. (17) with $\Gamma/\gamma=10^{6}$. The dotted line in Panel (a) represents the fidelity value $1-F=1.5\times 10^{-6}$, which is the expected value for the optimal choice of $\hat{O}$. The fidelity comes within $10^{-6}$ of this value after just one iteration and barely changes after that. The real and imaginary parts of the matrix element $O_{1,3}$ (and hence also $O_{3,1}$) shrink almost to zero after just one iteration, while all other matrix elements barely change in $10^{3}$ iterations, indicating that most matrix elements have almost no impact on the performance of the AQEC code.
Figure 4: Progression of the AQEC code with the optimization iteration number $k$ for 4-level system with uniform decay rates when all three components of the code are initialized to random values and subsequently optimized. The speed of convergence fluctuates as the optimization proceeds. Overall, the convergence speed is similar to that obtained when we optimized only the induced decay matrix $\hat{b}$.
@@ III.2 Qudit with uniform decay rates
If we take the code space $\left\{\left|1\right\rangle,\left|3\right\rangle\right\}$ and set $\Gamma/\gamma=10^{6}$ and $\gamma\tau=1$, we obtain the fidelity $F=0.9999985=1-1.5\times 10^{-6}$. This value is probably the best achievable value for the fidelity (at time $\tau=1/\gamma$). In fact, if we run the optimization algorithm using the above settings, the algorithm terminates immediately because of a computation error, which we believe is caused by the fact that the code that is provided as an initial guess is already at the maximum of the fidelity landscape. We note here that the fidelity can be expressed as $F=1-1.5\gamma/\Gamma$. Two thirds of the fidelity reduction, i.e. $\gamma/\Gamma$, can be attributed to the admixture of error states in the long-time dynamics. The rest, i.e. $\gamma/(2\Gamma)$, results from the relaxation at rate $\gamma^{2}/\Gamma$ under the influence of the AQEC protocol for duration $t=1/\gamma$. The factor of 2 is the same as the one in Eq. (13).
We now try the automated code search. In all calculations, we set $\gamma\tau=1$. To test the performance of the algorithm in optimizing the different components of the AQEC code, we perform a few different optimization tasks. First, we fix $\hat{O}=0$ and $\hat{b}$ as in Eq. (17) with $\Gamma/\gamma=10^{6}$, and we search for the optimal code space, starting the search from a randomly generated basis. The search converges to the code space $\left\{\left|1\right\rangle,\left|3\right\rangle\right\}$, and the fidelity comes within $10^{-8}$ of the maximum value, after only a few optimization iterations, as shown in Fig. 1.
Next we fix the code words to the states $\left\{\left|1\right\rangle,\left|3\right\rangle\right\}$ and $\hat{O}=0$, and we run the algorithm to optimize $\hat{b}$. The results of a typical run are shown in Fig. 2. As explained in Sec. II, $\hat{b}$ is initially filled with complex random numbers whose real and imaginary parts are each chosen from a uniform distribution in the range $[-0.5,0.5]$. The search typically reaches a fidelity of $F\approx 0.996$ after $10^{4}$ iterations and $F\approx 0.99969$ after $10^{5}$ iterations. The matrix $\hat{b}$ converges to a matrix where the matrix elements $b_{1,0}$ and $b_{3,2}$ are almost equal and have a large absolute value. After $10^{5}$ iterations, the absolute value of $b_{1,0}$ is typically on the order of $10^{2}$, the relative difference between $b_{1,0}$ and $b_{3,2}$ is on the order of $10^{-4}$, and the next largest matrix element is two orders of magnitude smaller than $|b_{1,0}|$. The slow convergence in this case can, at least partly, be explained by the fact that the optimal code described above requires having $b_{1,0}=b_{3,2}$ and $\left|b_{1,0}\right|\gg\gamma$, with the other matrix elements being equal to zero. The code does indeed evolve in this direction as the optimization procedure progresses. However, the inequality between $b_{1,0}$ and $b_{3,2}$ during the optimization creates a bottleneck for convergence. By considering the fidelity landscape, we can see that a small difference between $b_{1,0}$ and $b_{3,2}$ reduces the fidelity much more than any fidelity increase that can be gained by increasing the two values together by $|b_{1,0}-b_{3,2}|$. As a result, the algorithm identifies the optimal update policy as bringing the two values closer to each other, while slightly increasing the average. While the algorithm does converge in the correct direction, the approach to the optimal situation with the correct relations between the different matrix elements is rather inefficient and causes slow convergence.
Next, we fix the code words to the states $\left\{\left|1\right\rangle,\left|3\right\rangle\right\}$ and $\hat{b}$ as in Eq. (17) with $\Gamma/\gamma=10^{6}$. We optimize the operator $\hat{O}$, which we expect to have the optimal value $\hat{O}=0$. The results of a typical run are shown in Fig. 3. The fidelity consistently reaches a value between $0.999997$ and $0.999998$ in the first optimization iteration and remains constant (up to 12 significant figures) after only a few iterations. As a result, the search terminates as soon as the other early-termination conditions are satisfied, i.e. after about $10^{3}$ iterations. In most runs, the matrix elements $O_{1,3}$ and $O_{3,1}$ shrink to at most $\sim 1\times 10^{-6}$ in the optimized matrix $\hat{O}$, while all other matrix elements barely change from their initial random values, with net changes consistently below $10^{-4}$. The reason why only two matrix elements converge towards their optimal values is that all the other matrix elements involve states that remain almost unoccupied and therefore have a negligible effect on the fidelity. Only the matrix elements $O_{1,3}$ and $O_{3,1}$ cause a significant change in a general state in the code space. Even the matrix elements that mix states in the code space with states outside the code space (e.g. $O_{0,1}$) are effectively suppressed by the large induced decay described by the matrix $\hat{b}$. As is common in optimization problems, when a certain parameter has a negligible effect on the cost function (i.e. the fidelity in our case), the algorithm does not prioritize optimizing the parameter. For this reason, most matrix elements in $\hat{O}$ are barely updated and remain very close to their initial values. It is somewhat surprising that this stagnation in suppressing the almost irrelevant matrix elements occurs even when the fidelity is still $\sim 10^{-6}$ away from the maximum possible value, which is not an extremely small difference.
In about 10% of the calculations, the matrix elements $O_{1,3}$ and $O_{3,1}$ did not converge towards zero, but towards values that are integer multiples of $2\pi\gamma$ in magnitude. The reason is as follows: the dynamics induced by the matrix elements $O_{1,3}$ and $O_{3,1}$ is coherent and does not represent irreparable decoherence in the quantum state, provided that these matrix elements are known. Therefore, in principle, these matrix elements can take any value without contributing to decoherence. However, our optimization algorithm requires that the state remain at its initial value. Therefore, $O_{1,3}$ and $O_{3,1}$ must converge to values that take the code words back to their initial values when $\gamma\tau=1$, which is the time setting at which we evaluate the fidelity.
@@
Next, we do a full code search, in which we start with a randomly generated basis and randomly generated matrices $\hat{O}$ and $\hat{b}$. The fidelity for these random settings is typically in the range 0.25-0.5. Results from a typical run are shown in Fig. 4. In most runs, the fidelity reached values around $F=0.996$ after $10^{4}$ runs and $F=0.99965$ after $10^{5}$ runs, only slightly below the numbers that we obtained when we fixed the code space and optimized $\hat{b}$ only. We did not encounter any runs that seemed to get trapped in local optima with lower fidelities. All the components of the AQEC code converged towards the optimal values described above. By comparing this case with the the previously shown results, we conclude that the bottleneck for the fidelity convergence is the optimization of the induced decay matrix $\hat{b}$. Furthermore, the code basis and the coherent control matrix $\hat{O}$ both keep changing as long as $\hat{b}$ is changing, indicating that these components of the AQEC code keep finding optimal values that correspond to the current value of the matrix $\hat{b}$.
It is not difficult to find an intuitive generalization of the 13 code for an $n$-level system with $n>4$. In particular, if $n$ is even, the code space can be taken as $\left\{\left|n/2\right\rangle,\left|n\right\rangle\right\}$. Multiple induced-decay jump operators (which can also be combined into a single jump operator) are used to fix errors. Each one of these jump operators should fix an error associated with the loss of $k$ excitations, such that the quantum state returns to the code space by adding $k$ excitations, without affecting the quantum coherence between the states $\left|n/2-k\right\rangle$ and $\left|n-k\right\rangle$. One intuitive candidate for these jump operators is obtained by setting $b_{n/2-k,n/2}=b_{n-k,n}$ for $21\leq k\leq n/$ 2, with zero matrix elements otherwise. However, it turns out that this choice does not perform the desired task, because it merges states that had previously incurred different numbers of errors, which does not provide a mechanism to get rid of the entropy that is created in the system when the different errors occur. If instead we set $b_{n/2-k,n/2-k+1}=b_{n-k,n-k+1}$, we obtain a good AQEC code. For example, if we take $n=6$ and $\Gamma/\gamma=10^{6}$ for both $k=1$ and $k=2$, and we set $\gamma\tau=1$, we obtain a fidelity of $F=0.999999=1-1.0\times 10^{-6}$. The small reduction away from $F=1$ can now almost completely be accounted for in the admixture of error states in the long-time dynamics, with no noticeable contribution from the decoherence of the code words.
We ran the optimization algorithm on the case of a 6-level system. In the majority of the runs, the fidelity reached $F\approx 0.9992$ after $10^{5}$ iterations. The maximum value reached in one run (out of ten runs) was $F=0.99938$. These fidelity values are much higher than the free evolution value $F_{0}$ given in Sec. III.A, but short of the value 0.999999 obtained with intuitive reasoning above. The optimization algorithm is clearly able to find good AQEC protocols. We expect that it will converge to the optimal code if it is run longer. One caveat here is that there are a large number of extremely good AQEC codes in this 6-level system, such that the algorithm can already reach fidelities at the level of several nines with codes that are drastically different from the possibly optimal one described above. Furthermore, convergence tends to be slow when the fidelity approaches the maximum, mainly because of the slow convergence of the induced decay matrix $\hat{b}$. An additional complication that does not affect this particular example but can be important for other decay models is the fact that higher levels generally decay faster and result in higher admixtures of error states. Even if these admixtures do not affect the effective decay rate, they contribute a small reduction to the fidelity, making it difficult for the algorithm to distinguish between many different codes that are all extremely good.
III.3 Ququart with photon loss

### v2

@@ III.4 Optimization algorithm
The optimization algorithm is illustrated in Fig. 1. We start each instance of the search by making an initial guess for the AQEC code parameters. In most calculations, we make a random guess. As mentioned above, the AQEC code has three components: (1) the code space made up of the two code words, (2) the induced decay matrices, and (3) the controlled Hamiltonian matrix. The code space is chosen by taking two orthogonal but otherwise random vectors in the full Hilbert space. We use a single matrix $\hat{b}$ for the second component of the code. The matrices $\hat{b}$ and $\hat{O}$ are filled with complex numbers, such that each real or imaginary part of each matrix element is chosen from a uniform distribution in the range [0.5,0.5][-0.5,0.5]. One constraint is that $\hat{O}$ must be Hermitian. We therefore randomly generate the matrix elements above the diagonal and take their complex conjugates to calculate the matrix elements below the diagonal.
We use gradient-based techniques to optimize the three variable components of the AQEC code. We update each one of these three components by evaluating the gradient of the fidelity with respect to the relevant parameters and shifting the AQEC code parameters in the direction of the gradient. We alternate between updating the three components, such that each optimization iteration is composed of three separate update steps.
In principle, we could calculate a single, unified gradient that incorporates the parameters of all three components of the AQEC code. However, we chose to divide the step of updating the code parameters into three distinct update steps, alternating between the three and updating each one of them in each optimization iteration. There are a few reasons why we find the alternating update approach logical. For example, the code space updates are performed via applying unitary operations to the code space, while the matrices $\hat{b}$ and $\hat{O}$ are updated via simple addition. Furthermore, the optimal matrix $\hat{b}$ is expected to contain infinite matrix elements, as we will discuss in Sec. IV. We therefore expect some matrix elements to keep growing indefinitely but always remain infinitely far from the optimal values. These considerations make it conceptually simpler to keep the update steps separate. For comparison purposes, we repeated some of our calculations, specifically those presented in Sec. IV.B, using the unified-gradient approach where all three components are updated simultaneously. We found that these calculations take somewhat less time per optimization iteration but achieve a smaller fidelity improvement per iteration, such that the two approaches exhibit comparable overall convergence speeds.
@@
An optimization iteration proceeds as follows: Taking the best available code space, we consider that any neighboring basis can be obtained from the current basis using the update unitary operator
@@
where $\lambda_{m}$ are (small) coefficients and $G_{m}$ are the generators of unitary operators in the full Hilbert space, i.e. the generalized Gell-Mann matrices for the full Hilbert space. We generate a slightly modified basis for every generator, using the small infinitesimal coefficient $\lambda_{m}=10^{-8}$, and construct an array for the fidelity for the different possible updates. This array allows us to calculate the gradient to maximize the fidelity improvement. More specifically, we obtain a single generator that contains the optimal proportions of the different $G_{m}$ matrices. We can update the code words by applying to them the matrix $U_{\rm update}$ with the optimal generator multiplied by an update step size, which we initially set at $10^{-2}$. We keep moving along the direction of the gradient until the fidelity reaches a maximum and starts decreasing. When we identify that we have overshot the fidelity peak, we start zooming in on the peak region until we can identify the peak location to a precision of $10^{-8}$ in the update step size, i.e. until an update step of $10^{-8}$ in either direction decreases the fidelity rather than increases it. We update the code words to those that correspond to the point of maximum fidelity.
@@
For the controlled decay matrix $\hat{b}$, since it can be any $n\times n$ matrix, we take $2n^{2}$ possible updates for the real and imaginary parts of all matrix elements, each time adding a small increment to one of the matrix elements and calculating the fidelity with the small update. We use the increment size $10^{-6}$. Once we have the $2n^{2}$ values of fidelity for all the possible updates, we calculate the gradient and determine the optimal direction for updating $\hat{b}$. We then move in the optimal update direction until we reach a maximum in the fidelity. We note here that although we set a maximum for the allowed update in a single iteration ($10^{4}$ to any real or imaginary part of any matrix element), this condition was never needed by our numerical search algorithm. Including the condition was motivated by the fact that the optimal matrix $\hat{b}$ is expected to have some infinitely large matrix elements. When a peak in the fidelity is encountered during the update process, which is what happened in all iterations of all calculations, we start zooming in on the peak region until we can identify the maximum point location to within $10^{-8}$. A similar procedure is used to update the matrix $\hat{O}$, incorporating the constraint that the operator must be Hermitian.
@@
A natural question to consider here is whether the gradient-based approach is suitable for the problem of AQEC search, or whether this problem is prone to suffering from the presence of local optima that hinder the search for the optimal code. Our use of this approach was inspired by the results of Refs. Yamamoto2005 ; Yamamoto2007 ; Kosut , which alternated between optimizing the code space and the recovery operators in the conventional, measure-then-correct QEC paradigm. Those studies demonstrated that the search does not contain local minima, such that the gradient-based search and the alternation between updating different components of the code are good elements in a QEC code discovery algorithm. Our results in Sec. IV below show some indications that our search sometimes gets trapped in local optima and therefore fails to find optimal solutions. The authors of Ref. Wang made a similar observation and argued that the use of the average fidelity as an optimization cost function results in the appearance of local optima.
@@ IV Results
IV.1 Reference point: physical qubit with relaxation
To assess the AQEC protocols analyzed later in this section, we start by establishing the reference point against which the protocols should be evaluated.
@@ IV.1 Reference point: physical qubit with relaxation
The fidelity formula in Eq. (6) can then be evaluated by examining how much of each matrix element survives the relaxation dynamics. Concretely, we ignore the term $(1-e^{-\gamma t})\rho_{11}$ in the 00 matrix element, replace every $\rho_{ij}$ by 1, take the sum of all four matrix elements and divide by 4:
@@
One fidelity value that will serve as an important reference point below is the one obtained by setting $\gamma t=1$, because we use this setting in our numerical optimization calculations. With this setting, the fidelity under free evolution, i.e. when no error-correction action is taken, is $F_{0}=0.64523519$.
@@
IV.2 Ququart with uniform decay rates
Figure 2: Progression of AQEC code with optimization iteration number $k$ for 4-level system with uniform decay rates. The infidelity (1F)(1-F), plotted in Panel (a), decreases steadily and approaches zero as a function of $k$, indicating that the algorithm is successful in finding a good AQEC code. Panel (b) shows that the code space approaches $\left\{\left|1\right\rangle,\left|3\right\rangle\right\}$. In Panel (c) the legend specifies only the highest two lines. In this specific run, by coincidence, the real parts of $b_{1,0}$ and $b_{3,2}$ did not grow significantly up to $k=10^{5}$. In Panel (d), the inset shows the real (cyan) and imaginary (gray) parts of $O_{1,3}$. This matrix element is expected to converge to zero in the optimal code. Most other matrix elements in $\hat{O}$ are expected to have a small effect on the fidelity, as explained in Appendix A. All plots in this figure indicate that the algorithm is progressing towards the 13 code but that it has not converged yet.
@@ IV.2 Ququart with uniform decay rates
If we take the code space $\left\{\left|1\right\rangle,\left|3\right\rangle\right\}$ and set $\Gamma/\gamma=10^{6}$ and $\gamma\tau=1$, we obtain the fidelity $F=0.9999985=1-1.5\times 10^{-6}$. This value is probably the best achievable value for the fidelity (at time $\tau=1/\gamma$). In fact, if we run the optimization algorithm using the above settings, the algorithm terminates immediately because of a computation error, which we believe is caused by the fact that the code that is provided as an initial guess is already at the maximum of the fidelity landscape. We note here that the fidelity can be expressed as $F=1-1.5\gamma/\Gamma$. Two thirds of the fidelity reduction, i.e. $\gamma/\Gamma$, can be attributed to the admixture of error states in the long-time dynamics. The rest, i.e. $\gamma/(2\Gamma)$, results from the relaxation at rate $\gamma^{2}/\Gamma$ under the influence of the AQEC protocol for duration $t=1/\gamma$. The factor of 2 here is the same as the one in Eq. (13).
@@
We now try the automated code search. As mentioned above, we start with a randomly generated basis and randomly generated matrices $\hat{b}$ and $\hat{O}$, and we set $\gamma\tau=1$. The fidelity for these random settings is typically in the range 0.25-0.5. Results from a typical run of the optimization algorithm are shown in Fig. 2. In most runs, the fidelity reached values around $F=0.996$ after $10^{4}$ iterations and $F=0.99965$ after $10^{5}$ iterations. We did not encounter any runs that seemed to get trapped in local optima with lower fidelities. All the components of the AQEC code converged towards the optimal values described above.
In Appendix A, we present the results of optimizing the three AQEC code components individually. By comparing the results of this subsection with those of Appendix A, we conclude that the bottleneck for the fidelity convergence is the optimization of the induced decay matrix $\hat{b}$. Furthermore, the code basis and the coherent control matrix $\hat{O}$ both keep changing as long as $\hat{b}$ is changing, indicating that these components of the AQEC code keep finding optimal values that correspond to the current value of the matrix $\hat{b}$.
It is worth making a few comments about the model treated in this section before moving on to the next model. The four-level model could be seen as a toy model that is useful as a tool to gain a basic understanding of the idea of AQEC. At the same time, it is quite conceivable that a physical system, e.g. a superconducting circuit with a cleverly designed arrangement of Josephson junctions, can have a quantum state structure with transition rates that allow the realization of the 13 code. It is important to note here that for the 13 code to work, the $\left|2\right\rangle\to\left|1\right\rangle$ transition rate does not need to have any particular relation to the $\left|1\right\rangle\to\left|0\right\rangle$ and $\left|3\right\rangle\to\left|2\right\rangle$ transition rates. The only requirement is that the latter two be equal and that these transitions not leave a signature of the exact pair involved in the transition, e.g. by emitting photons of different frequencies.
IV.3 Ququart with photon loss